# Efficient Cloud-Edge Collaborative Inference for Object Re-Identification

### v1

@@ 5.4 Ablation Study
Figure 4: The effects of different values of $\gamma_{0}$ (left) and $\gamma_{1}$ (right) for mTN.
Figure 5: The system performance with different bandwidth limits.
Influence of different bandwidth limits $B$. The bandwidth limits $B$ are important for the ReID system, thus we explore its effect by setting it to 16, 24, 32, respectively, and see the changing of mTN for different strategies. The results are shown in Fig. 5, and we select Average, Ratio, stReID, and ours as the comparison strategies. Their details can be referred to the supplementary materials. From the results, we can learn that in all three Settings, our approach achieves the lowest number of data transfers.
Figure 6: The visualization of retrieval results for some samples from the DukeMTMC-reID dataset. Two sections for baseline and our approach are shown. For each section, the left part represents the query, followed by the ranked retrieval results from the gallery. Persons distinct from the query are highlighted with a red border.
@@ 5.5 Visualization
To help understand the effect of our approach, we visualized some examples that are corrected by the learned spatial-temporal correlations. The same examples are shown in Fig. 6. The red box indicates inconsistency with the query image ID, while the green box indicates consistency. It is evident that utilizing correlation results in the retrieval of more correct images, leading to the improvement in mAP. More visualized examples please refer to our supplementary materials.
@@ 6 Conclusion
In conclusion, the increasing volume of video content makes the traditional centralized ReID system impractical. The current cloud-edge collaborative methods face challenges related to bandwidth constraints and search efficiency. To address these problems, our study introduces a pioneering cloud-edge collaborative ReID framework. Leveraging a distribution-aware correlation modeling network, our approach enables efficient inference, ensuring the desired image returns to the cloud server as early as possible. Comparative experiments with existing ReID methods demonstrate our approach can reduce a huge amount of data transmission, improve performance, and achieve time-constraints ReID, showing its superiority. We also find that the proposed approach requires that the data must have high-quality timestamps, otherwise, the performance will be seriously degraded, which is why we mainly conduct experiments on the DukeMTMC-reid dataset, while other datasets often have various problems with timestamps. Thus, how to use low-quality time information to achieve high-quality correlation learning may be our future research direction.

### v2

@@ 5.4 Ablation Study
Figure 6: The visualization of retrieval results for some samples from the DukeMTMC-reID dataset. Timestamps (Frame ID) are shown below the images.
@@ 5.5 Visualization
To help understand our approach, we provide some visualization examples to illustrate the impact of our proposed method, as depicted in Fig. 6. Instances marked with a red box signify inconsistency with the query image ID, while those marked with a green box indicate consistency.
These visualizations underscore the challenges of distinguishing certain images based solely on visual appearance. For instance, in the initial set of images (a), the individual in the third image is in a blue long-sleeved shirt and black pants, sharing a notable resemblance with the person in the query image. In the second row of the visualization results, the woman in the first two images of (d) wears a black hoodie, denim pants, and black boots, and holds white rolls of paper, exhibiting a noticeable similarity in appearance with the person in the query image. However, as can be seen from (f), the pedestrian in the correct images is facing away from the camera, and the coat does not exhibit any features of the white paper roll. In contrast, the person in the third image of (e) is attired in all black, and lacks a hat, but shares a dark hair color, bearing a strong resemblance to the pedestrians depicted in the two returned images. However, the stReID method fails to filter out this incorrect result based on the spatial-temporal statistic approach, and our method effectively filters out unreliable returned images.
@@ 6 Conclusion
The increasing volume of videos makes the traditional centralized ReID system impractical, and the current cloud-edge collaborative methods face challenges related to bandwidth constraints and search efficiency. To address these problems, we introduce a pioneering cloud-edge collaborative ReID framework. By leveraging a distribution-aware correlation modeling network, our approach enables efficient inference, ensuring the desired image returns to the cloud server as early as possible. Comparative experiments demonstrate our approach can reduce data transmission and improve the performance across various baselines, showing its superiority. We also acknowledge that it requires time stamps and is inappropriate for mobile devices, which motivate our future research on utilizing unstable spatial-temporal data to achieve high-quality correlation learning.
Acknowledgments
This work was supported in part by the Funds for the NSFC Project under Grant U24B20176, 62202063, 62406038, 62302058, and China Postdoctoral Science Foundation under Grant 2024M760280.